# Selective labeling with false discovery rate control

### v1

@@ 5 Related work
Selection with conformal $p$-values.
Conformal $p$-values are widely studied for their distribution-free, model-agnostic properties in selection tasks. Existing methods fall into two main categories: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025; Huo et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024). Conformal novelty detection aims to identify out-of-distribution instances, while conformal selection aims to select data points that meet a specific quality criterion. For conformal novelty detection, recent advances primarily focus on improving the power of the selection procedure by using various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025). Jin & Candès (2023) introduced the conformal selection framework for data selection in regression settings, which has since been extended to various settings: multivariate data selection (Bai et al., 2025b), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025). Recent works have also applied CS to different tasks such as candidate screening (Lu et al., 2025), drug discovery (Jin & Candès, 2023; Bai et al., 2025a), and foundation model alignment (Gui et al., 2024). However, none of them has studied how to construct conformal $p$-values and design selection procedures in the selective labeling task.
Selective labeling.
Our method builds on the idea of selective labeling, which prioritizes the collection of expert labels for instances where the AI model exhibits uncertainty, and relies on the model’s prediction where it is confident. (Gu et al., 2012; Vrabac et al., 2022). Prior works on selective labeling have explored various heuristic methods, including collaborative annotation frameworks that integrate expert and LLM efforts (Li et al., 2023), and domain-specific applications in text, vision, and medicine (Kim et al., 2024; Vrabac et al., 2022; Duan & Lalor, 2023; Zhang et al., 2025). More recent research aims to provide theoretical foundations, with methods designed to control the overall labeling error (Candès et al., 2025) or to enable valid statistical inference (Zrnic & Candes, 2024; Gligorić et al., 2024). Our method is also closely related to selective prediction, where models are allowed to abstain from making a prediction when uncertain about the output (Geifman & El-Yaniv, 2017; Mozannar & Sontag, 2020; Yang et al., 2023). For LLMs, recent works have also studied how to teach the model not to predict when the model is uncertain (Kamath et al., 2020; Yoshikawa & Okazaki, 2023). These selective prediction methods also lack theoretical guarantees for the AI predictions. Distinct from prior methods, our work is the first to provide a guarantee for the quality of AI-assigned labels, thereby enabling the reliable deployment of selective labeling.
@@ 6 Conclusion
In this work, we propose Conformal Labeling, a novel selective labeling method for identifying samples where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), which ensures that the expected fraction of incorrect labels in the selected subset is below a user-specified level. The key idea is to reformulate selective labeling as multiple hypothesis testing, which enables distinct theoretical guarantees and methodological advantages compared to prior approaches. In particular, we construct a conformal $p$-value for each test instance by comparing the AI model’s predicted confidence to those of mislabeled calibration instances. Then, we select all the test samples whose $p$-values are smaller than or equal to a data-dependent threshold. We theoretically prove that Conformal Labeling successfully controls the FDR under mild assumptions. Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image and text labeling, and LLM QA. We hope the insight from this work will inspire future research to explore reliable selective labeling.
Appendix A Use of Large language model
This paper uses large language models solely to polish specific sentences or paragraphs, without further use of LLMs for other purposes.
Appendix B Technical proofs
B.1 A lemma for proving Theorem 3.1
Lemma B.1.
@@ Appendix B Technical proofs
Suppose the first $n_{0}$ are true null samples in the calibration set and the first $m_{0}$ samples are true null samples in the test set. By the standard result in conformal inference (Vovk et al., 2005), we have
$\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid n_{0},m_{0}\}\leq\alpha,j=1,\ldots,m_{0}$
Let $j^{*}$ denote the number of rejections from Algorithm 1, and recall the false discovery proportion (FDP) as
$\text{FDP}=\frac{\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}}{\max(j^{*},1)}.$
Since $\text{FDP}=0$ whenever $j^{*}=0$, we have
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]=$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]\mathbb{P}\left\{j^{*}>0\right\}$ (3)
$\quad\quad+\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}=0\right]\mathbb{P}\left\{j^{*}=0\right\}$
$\leq$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]$
$=$ $\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right].$
By the super-uniform property of conformal $p$-values,
$\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\;\middle|\;n_{0},m_{0},j^{*}\right]\leq m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}.$
Thus,
$\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{1}{j^{*}}\cdot m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}=\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha.$
Combining with (3) yields
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha,$
which completes the proof. ∎
B.2 Proof of theorem 3.1
@@
Appendix C Overview of Different Uncertainty Score Functions
@@
Appendix D BH Procedure and its Adaptive Variants
@@
If $\pi_{0}$ is small, the FDR control will be overly conservative. When $\pi_{0}$ is known, we can apply BH procedure at level $\frac{\alpha}{\pi_{0}}$ to close the gap. In practice, $\pi_{0}$ is typically unknown. Several adaptive BH procedures attempt to address this issue by estimating $\pi_{0}$ and adjusting the target FDR level $\alpha$ accordingly. These procedures are often called the $\pi_{0}$-adaptive versions of the BH algorithm. Two most famous estimators are Storey-BH (Storey, 2002) and Quantile-BH (Benjamini et al., 2006):
$\hat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
$\hat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
@@
Appendix E Hyperparameter selection for BH adaptive variants
@@
    Select $\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
@@ Appendix F Implementation details
Experiment details. We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and vLLM.
Dataset details.
For the LLM QA datasets (MedMCQA, MMLU, and MMLU-Pro), we adopt the same prompts as in Luo et al. (2025). For the other datasets, we design our own prompts following the style of LLM QA prompts, as summarized in Table 2. In text labeling and LLM QA tasks, we merge the calibration and test sets to form the unlabeled dataset, except for MedMCQA. For MedMCQA, because test labels are unavailable, we use the calibration set as the unlabeled dataset. For image labeling, we use the validation sets of ImageNet and ImageNet-V2 as the unlabeled datasets.
Table 2: Prompts for different datasets
   Dataset    Prompts    LLM QA    The following are multiple-choice questions. Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]        Stance on    global warming    You are given a statement about climate change. Determine the stance that a human would take towards this statement.    Respond with ONLY the letter (A, B, or C) of the correct stance. Do not include any explanation.    [Input Headline] A: agrees B: neutral C: disagrees Answer: [Mask]    Misinformation    You are a fact-checking assistant. Classify the following news headline as either real (A) or misinfo (B).    Respond with ONLY the letter A or B. Do not include any explanation.    Headline: [Input text] A: real B: misinfo Answer: [Mask]
Appendix G Extension to Regression Tasks
@@
While our primary focus is on classification, Conformal Labeling can be naturally extended to regression settings. Consider a loss function L(Y,Y^)L($\hat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\hat{Y})=(Y-\hat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$. For each test sample, we define the null and alternative hypotheses as:
$H_{0}^{j}:L(Y_{n+j},\hat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\hat{Y}_{n+j})\leq\epsilon.$
@@
This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\hat{Y})=\mathbf{1}\{Y\neq\hat{Y}\}$ and $\epsilon=0$. To apply Conformal Labeling in regression, we need an uncertainty score that reflects the model’s predictive uncertainty. While uncertainty score functions are straightforward in classification tasks (e.g., $1-\max_{y\in\mathcal{Y}}f_{y}(X)$), regression requires alternative approaches to get an uncertainty score function. For example, when using LLMs, we can leverage their verbalized confidence or prompt them to output prediction intervals, using the interval width as the uncertainty score. The complete procedure for Conformal Labeling in the regression task is outlined in Algorithm 2.
Algorithm 2 Conformal Labeling Regression Tasks
@@
1:Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
2:Identify calibration samples that exceed the loss threshold: $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\hat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and set $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
3:for $j=1,\dots,m$ do
4:  Compute the conformal $p$-value $\hat{p}_{j}$ according to equation 2.
5:end for
@@
6:Apply the step-up procedure: $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest $p$-value.
7:Output: The selected set $\mathcal{S}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
@@
We evaluate Conformal Labeling on two regression problems: sentiment analysis with GPT-4o and protein structure prediction with AlphaFold. We use the data provided by Candès et al. (2025). For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\hat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score. For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\hat{Y}_{i}$. Uncertainty scores are obtained from AlphaFold’s internal confidence measure, the average predicted local distance difference test (pLDDT). For both experiments, we employ the L2 loss function.
In Table 3, we report the performance of Conformal Labeling on regression tasks with $\alpha=0.1$. In all cases, Conformal Labeling consistently controls the realized FDR below the target level. A key observation is that Conformal Labeling’s selection result in regression task is highly sensitive to the choice of the tolerance parameter $\epsilon$. For instance, in sentiment analysis, setting $\epsilon=0.05$ yields an AI-labeled ratio of only $4.55\%$, whereas a slightly larger value $\epsilon=0.06$ increases the ratio to $47.84\%$. In practice, selecting an appropriate $\epsilon$ may require domain expertise or prior knowledge about the task.

### v2

@@ 5 Related work
Conformal Inference.
Conformal inference (Vovk et al., 2005; Gazin et al., 2025; Humbert et al., 2025) is a statistical framework characterized by distribution-free and model-agnostic theoretical guarantees. In the literature, two particularly popular directions have emerged: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024). The former focuses on identifying out-of-distribution instances, with recent advances focusing on enhancing selection power by incorporating various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025). The latter aims to select candidates whose unobserved outcomes exceed user-specified values (Jin & Candès, 2023), which has been extended to multivariate data selection (Bai et al., 2025), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025). Besides, some works have developed conformal inference techniques for classification tasks (Zhao & Su, 2023; Sun & Xia, 2025). For example, recent work (Zhao & Su, 2023) proposes controlling the overall error rate in binary and multi-class classification, while a concurrent work (Sun & Xia, 2025) extends this to control the general group-wise false discovery rate within a unified framework. In this work, we motivate and tailor the framework of conformal selection for selective labeling and support the application for open-ended generation tasks.
Selective labeling and prediction.
Selective labeling is an increasingly important topic for balancing the tradeoff between labeling cost and error (Gu et al., 2012; Vrabac et al., 2022). While prior work mainly explored heuristic approaches (Li et al., 2023; Duan & Lalor, 2023), a concurrent work (Candès et al., 2025) proposes approximately correct labeling. It guarantees that the overall labeling error is small with high probability by counterbalancing the error of AI labels with the zero error of expert labels. Our work considers a different objective: controlling the labeling error within the AI-labeled subset. Thus, these two methods address different error notions and are applicable in different settings. Notably, our method does not require expert annotations for samples on which the AI model abstains, whereas PAC labeling leverages such annotations to control the overall error. Some other work has also taken steps toward providing statistical guarantees for selective labeling (Zrnic & Candes, 2024; Gligorić et al., 2024), though these works focus on downstream inference rather than guaranteeing label quality. Our method is also related to selective prediction, where a model is allowed to abstain from making predictions when uncertain (El-Yaniv & Wiener, 2010; Mozannar & Sontag, 2020; Yang et al., 2023; Kamath et al., 2020; Yoshikawa & Okazaki, 2023). However, most of these methods cannot provide theoretical guarantees on prediction error. While works on selective prediction with risk control (El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017) can provide theoretical guarantees, they target a different error measure called marginal FDR, which is less preferred by prior researchers compared with FDR control. We provide a detailed discussion of the error measures in Appendix I.4.
@@ 6 Conclusion
In this paper, we introduce Conformal Labeling, a novel method for identifying a subset where AI predictions can be provably trusted. Specifically, we construct conformal $p$-values and estimate the accuracy of AI models using a small labeled calibration dataset. Then, we design a principled selection procedure that automatically adapts to the estimated accuracy, with higher accuracy permitting a larger threshold. We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level under mild assumptions. Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image labeling, LLM QA, and LLM open-ended generation tasks. We hope that the insights from this work will inspire future research on rigorous methods for selective labeling.
Appendix A Technical proofs
A.1 Lemmas for proving Theorem 3.2
Lemma A.1.
(Benjamini & Yekutieli (2001), Theorem 1.2) If the joint distribution of the $p$-values is PRDS on the subset corresponding to true null hypotheses, applying the Benjamini–Hochberg procedure at level $\alpha$ guarantees
$\text{FDR}\leq\frac{m_{0}}{m}\alpha,$
where $m_{0}$ is the number of true null hypotheses.
Lemma A.2.
@@ Appendix A Technical proofs
Without loss of generality, assume that the first $n_{0}$ calibration samples and the first $m_{0}$ test samples correspond to true null hypotheses. By the standard result in conformal inference (Vovk et al., 2005) we have,
[…]
From results in (Bates et al., 2023), we know conformal $p$-values $(\widehat{p}_{1},\ldots,\widehat{p}_{m})$ are PRDS on the set of mislabeled data. Lemma A.1 hence guarantees that conditional on $n_{0}$ and $m_{0}$, applying Conformal Labeling at level $\alpha$ satisfies
A.2 Proof of theorem 3.2
@@
Appendix B Overview of Different Uncertainty Score Functions
@@
Appendix C BH Procedure and its Adaptive Variants
@@
If $\pi_{0}$ is small, the FDR control will be overly conservative. For example, if a classifier achieves a 80% accuracy in the unlabeled test dataset, then $\pi_{0}=0.2$—which leads to overly conservative FDR control. When $\pi_{0}$ is known, we can apply BH procedure at level $\frac{\alpha}{\pi_{0}}$ to close the gap. In practice, $\pi_{0}$ is typically unknown. Several adaptive BH procedures attempt to address this issue by estimating $\pi_{0}$ and adjusting the target FDR level $\alpha$ accordingly. These procedures are often called the $\pi_{0}$-adaptive versions of the BH algorithm. Two most famous estimators are Storey-BH (Storey, 2002) and Quantile-BH (Benjamini et al., 2006):
$\widehat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
$\widehat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
@@
Appendix D Hyperparameter selection for BH adaptive variants
@@
    Select $\widehat{\lambda}=\operatorname*{argmin}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
Appendix E Uncertainty score functions for open-ended generation task
Logits-based uncertainty score.
In particular, let $y=(y_{1},y_{2},\ldots,y_{l})$ be a generated sequence of length $l$. We define the logits-based score as:
$S_{\text{logits}}(x)=1-\frac{1}{l}\sum_{j=1}^{l}\mathbb{P}(y_{j}\mid y_{1},\ldots,y_{j-1},x)$
where $\mathbb{P}(y_{j}\mid y_{1},\ldots,y_{j-1},x)$ is the conditional probability of token $y_{j}$. This score measures the average confidence of the model across all generated tokens. When the model is very confident about each token, the score is low. When the model is less confident, the score is high, indicating an example that may benefit from a more capable annotation source.
Verbalized uncertainty score
For the verbalized score, we first let the model explicitly state its self-reported confidence. The uncertainty score is then defined as one minus the self-reported confidence. In this work, we report the average confidence over 10 trials, and the prompts are listed in Table 4.
Table 4: Prompt for the verbalized confidence scores.
    System prompt: You are a reasoning assistant. For each question and proposed answer, you must estimate how likely the proposed answer is correct.
    User prompt:
    Question: {QUESTION}
    Answer: {ANSWER}
    Provide a probability (between 0.0 and 1.0) that your answer is correct. Only output the probability.
@@ Appendix F Implementation details
Hyperparameter settings of LLMs
For the open-ended generation task, we configure the decoding parameters as follows. For Qwen/Qwen3-4B-Instruct-2507, we set Temperature=0.7, TopP=0.8, TopK=20, and MinP=0. For Qwen/Qwen3-4B-Thinking-2507, we set Temperature=0.6, TopP=0.95, TopK=20, and MinP=0. For DeepSeek/DeepSeek-R1-Distill-Qwen-32B, we set Temperature=0.6, TopP=0.95, and MinP=0. All experiments were run on NVIDIA GeForce RTX 4090, NVIDIA L40, and NVIDIA RTX Pro 6000.
Dataset splitting strategy.
For image labeling and LLM QA tasks, we randomly split 10\$ of the unlabeled dataset as the calibration dataset. For open-ended generation task, we use dataset-specific splitting strategy, which is detailed in Table LABEL:.
Table 5: Details of dataset splitting settings for open-ended generation task.
Dataset Dataset Type Dataset Size Split Setting Size
MMLU-Redux General Task 3000 Calibration 500
Test 2500
MATH-500 Math Reasoning 500 Calibration 300
Test 200
MATH-L5 Math Reasoning 721 Calibration 300
Test 421
Zebra-Logic Text Reasoning 750 Calibration 300
Test 450
HumanEval+ Coding Task 164 Calibration 100
Test 64
Prompts for LLM QA task
We adopt the same prompts as in Luo et al. (2025). The prompt is presented in Table 6.
Table 6: Prompt for LLM QA task
   Dataset    Prompt    LLM QA    The following are multiple-choice questions. Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]
Appendix G Compared baselines
We compare Conformal Labeling against four baseline methods: (1) an FDR-search heuristic baseline; (2) selection with guaranteed risk control (SGR) (Geifman & El-Yaniv, 2017); (3) labeling test instances with AI models when the uncertainty score satisfies $\mathcal{S}_{n+j}\leq\alpha$; and (4) applying AI predictions to the entire test dataset. FDR-search heuristic baseline selects the largest uncertainty threshold such that the empirical FDR of the calibration dataset does not surpass $\alpha$. SGR is a theoretically rigorous method that provides high-probability guarantees on marginal FDR control. It searches for the cutoff over a grid of confidence values, and then takes a union bound to achieve the overall error control. We provide a detailed introduction to the FDR-search heuristic baseline and SGR below.
FDR-search heuristic baseline.
Let $S(X)$ be a confidence function such as MSP. For a confidence threshold $t$, we define the empirical FDR on the calibration set as
$\widehat{\mathrm{FDR}}(t)=\frac{\sum_{i=1}^{n}\mathbf{1}\{S(x_{i})\geq t,\;Y_{i}\neq\widehat{Y}_{i}\}}{\max\left\{\sum_{i=1}^{n}\mathbf{1}\{S(x_{i})\geq t\},\;1\right\}}$
The threshold is selected by searching from the largest to the smallest confidence value. Specifically, we start from the smallest possible threshold and progressively increase $t$ until the empirical FDR on the calibration set falls below the target level $\alpha$:
$t^{\ast}(\alpha)=\min\{\,t:\widehat{\mathrm{FDR}}_{\mathrm{cal}}(t)\leq\alpha\,\}.$
Then a test instance is selected if its uncertainty score satisfies $\mathcal{S}(X_{i})\geq t^{\ast}(\alpha)$.
Selection with guaranteed risk control (SGR).
Let $f$ be a fixed classifier and let $c:\mathcal{X}\to\mathbb{R}$ be a confidence score, where larger values indicate more reliable predictions. For a threshold $\theta$, define the selection rule
$g_{\theta}(x)=\mathbf{1}\{c(x)\geq\theta\},$
so that predictions are made only on instances with confidence at least $\theta$. For a calibration set $\{(x_{i},y_{i})\}_{i=1}^{n}$, the empirical selective risk at threshold $\theta$ is
$\widehat{R}(\theta)=\frac{\sum_{i=1}^{n}\mathbf{1}\{g_{\theta}(x_{i})=1,\;f(x_{i})\neq y_{i}\}}{\max}\left\{\sum_{i=1}^{n}\mathbf{1}\{g_{\theta}(x_{i})=1\},\,1\right\}.$
The SGR algorithm searches over candidate thresholds (ordered from high to low confidence) and, for each threshold, computes a high-probability upper bound on the true selective risk based on the calibration data. The output threshold $\theta^{\ast}$ is the largest threshold whose risk bound does not exceed a target level $r^{\ast}$. At test time, an instance is selected if $c(X_{i})\geq\theta^{\ast}$.
Theorem G.1 (Selective risk guarantee of SGR).
Fix a target risk level $\alpha\in(0,1)$ and a confidence parameter $\delta\in(0,1)$. Let $\theta^{\ast}$ be the threshold returned by the SGR algorithm using a calibration set of size $n$. Then, with probability at least $1-\delta$ over the draw of the calibration set, the resulting selective classifier $(f,g_{\theta^{\ast}})$ satisfies
$R(f,g_{\theta^{\ast}})\;\leq\;\alpha,$
where $R(f,g_{\theta^{\ast}})$ denotes the true selective risk.
In the labeling task, the risk function is defined as
$R=\frac{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{Y_{n+j}\neq\widehat{Y}_{n+j}\text{ and }j\in\mathcal{R}\}\right]}{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{j\in\mathcal{R}\}\right]},$
which is also known as the marginal false discovery rate (mFDR) in hypothesis testing. In the literature, FDR control is generally preferred over mFDR control. We provide a discussion of mFDR and FDR control in Appendix I.4.
Appendix H Extension to Regression Tasks
@@
While our primary focus is on classification, Conformal Labeling can be naturally extended to regression settings. Consider a loss function L(Y,Y^)L($\widehat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\widehat{Y})=(Y-\widehat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$. For each test sample, we define the null and alternative hypotheses as:
$H_{0}^{j}:L(Y_{n+j},\widehat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\widehat{Y}_{n+j})\leq\epsilon.$
@@
This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\widehat{Y})=\mathbf{1}\{Y\neq\widehat{Y}\}$ and $\epsilon=0$. To apply Conformal Labeling in regression, we need an uncertainty score that reflects the model’s predictive uncertainty. While uncertainty score functions are straightforward in classification tasks (e.g., $1-\max_{y\in\mathcal{Y}}f_{y}(X)$), regression requires alternative approaches to get an uncertainty score function. For example, when using LLMs, we can leverage their verbalized confidence or prompt them to output prediction intervals, using the interval width as the uncertainty score. The complete procedure for Conformal Labeling in the regression task is outlined in Algorithm 2.
Algorithm 2 Conformal Labeling for Regression Tasks
@@
0: Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
0: Selected set $\mathcal{S}$ with FDR control
1:// Step 0: Identify calibration samples exceeding the loss threshold
2: Set $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\widehat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and let $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
3:// Step 1: Construct conformal $p$-values
4:for $j=1$ to $m$ do
5:  Compute the conformal $p$-value $\widehat{p}_{j}$ according to Eq. (3).
6:end for
7:// Step 2: Step-up thresholding
@@
8: Compute $j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\widehat{p}_{(j)}$ is the $j$-th smallest $p$-value.
9:Return $\mathcal{S}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$.
@@
We evaluate Conformal Labeling on two regression problems: sentiment analysis with GPT-4o and protein structure prediction with AlphaFold. We use the data provided by (Candès et al., 2025). For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\widehat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score. For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\widehat{Y}_{i}$. Uncertainty scores are obtained from AlphaFold’s internal confidence measure, the average predicted local distance difference test (pLDDT). For both experiments, we employ the L2 loss function.
In Table 7, we report the performance of Conformal Labeling on regression tasks with $\alpha=0.1$. In all cases, Conformal Labeling consistently controls the realized FDR below the target level, while a larger loss threshold $\epsilon$ leads to higher power.